OpenStudy (handymani):

Solve the following system of linear equations by substitution. x + 3y = 5 4x + 5y = 13

OpenStudy (will.h):

Oh they already said substitution. So okay We will have to set one of the equations with one variable isolated on one side

OpenStudy (will.h):

Can you do this?

OpenStudy (seratul):

x + 3y = 5 can be written as x = 5 + 3y Try substituting x in the next equation for 5 + 3y. 4(5 + 3y)+ 5y = 13

OpenStudy (will.h):

Agreed. And then simplify